OpenStudy (anonymous):

Trevor tutors French for $15 and hour and scoops ice cream for $10 an hour. He is going to work 15 hours this week

OpenStudy (abhisar):

Hello @lateriusb ! Do you think your question is complete?

OpenStudy (abhisar):

What I believe that your question is something like this "Trevor tutors French for $15 an hour and scoops ice cream for $10 an hour. He is going to work 15 hours this week. What will be his total earning?" Is it correct?

OpenStudy (abhisar):

Kindly clarify, we can not help you with incomplete questions. Thanks.

OpenStudy (anonymous):

Having internet trouble. I will resend it

OpenStudy (abhisar):

ok

OpenStudy (abhisar):

@lateriusb Since you never clarified your actual question so I am intended to explain another similar question which I think is close to your actual question. It may help you and other looking for similar question "Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?"

OpenStudy (abhisar):

Now, lets try to understand it part by part. For Part A: Let's see what's already given in the question. Number of working hours per week = \( x\) Money Drey gets for working each normal hour = $10 Wage earned = W So, we can draft an equation as W= 10\(x\)

OpenStudy (abhisar):

Similarly for Part B: What's given: Total wage = S Number of overtime hours he worked = y Money he gets for each overtime hour = 15$ Now, one thing you should keep in mind that for doing any hour of overtime Drey should do minimum of 40 hours of normal work. So the money he will be paid for doing y hours of overtime will be 40 hours of normal work with a rate of 10$ + y hours of overtime with an extra rate of 15$. Hence, we can draft an equation as follows, 40\(\times 10$\) + 15$\(\times\) y = S

OpenStudy (abhisar):

For Part C, try to do yourself, fell free to reach me if you have any doubt or problem. Thanks.